# SoK: Payment Channel Networks

### v2

@@ Abstract
Payment Channel Networks (PCNs) have been proposed as an alternative solution to the scalability, throughput, and cost overhead associated with on-chain transactions. By facilitating offchain execution of transactions, PCNs significantly reduce the burden on the blockchain, leading to faster transaction processing, reduced transaction fees, and enhanced privacy. Despite these advantages, the current research in PCNs presents a variety of research challenges that require further exploration.
In this paper, we survey the recent work in several aspects of PCNs, such as pathfinding and routing, virtual channels, state channels, payment channel hubs and rebalancing. This survey aims to provide the reader with a detailed understanding of the current state-of-the-art in PCN research, highlighting a few important advancements. Additionally, we highlight the various unresolved issues in the area of PCN research.
Specifically, this paper seeks to answer the following crucial question: What are the various interesting and non-trivial challenges in PCN research that require immediate attention from the academic and research community? By addressing this question, we aim to identify the most pressing problems and future research directions that interested readers can immediately work on. Through this analysis, we hope to inspire researchers and practitioners to tackle these challenges to make PCNs more secure and versatile.
@@ I Introduction
Cryptocurrencies and cryptocurrency based transactions have become increasing popular. Currently, the total market value of all cryptocurrencies in use has surpassed 2.5 Trillion USD. The cyrptocurrency market is increasing at a rate of $\approx$ 8.00% every year [1]. This rise in popularity can be attributed to the following reasons: 1) cryptocurrency transactions can be carried without the presence of a trusted entity. Fiat currency based transactions on the other hand, require the presence of a trusted financial organization such as a bank. 2) cryptocurrency transactions do not subject the user to any limits on the number and type of transactions. Fiat currency based transactions are limited in their amount and number, depending on several factors such as the currency, geographical location, etc. [2].
Each transaction posted to the Bitcoin blockchain takes around 7 seconds to be validated [3, 4]. The procedure of validation involves verifying that the transaction posted to the blockchain contains all the required fields and the signature of the user creating the transaction tuple is valid. Once the validation procedure is successfully completed, the transaction is included in a block that would be mined on the blockchain at some time in the future. The process of mining the block successfully takes $\approx$ 2 hours [5] (as of June $13^{th}$ 2024). This delay in the transactions and blocks getting confirmed is termed as the blockchain scalability problem [6]. In contrast, Visa, a company which globally processes transactions using fiat currencies, processes around 24,000 transactions per second [7]. Due to the delay in transaction processing caused by the blockchain scalability problem, blockchain-based transactions cannot process payments instantaneously.
@@
The idea of a payment channel that exists between two parties can be extended to a number of nodes, creating a network of payment channels, called a payment channel networks or PCN. The advantage of PCNs is, they enable users that are not connected by a direct payment channel to make payments between each other in an off-chain manner. An example PCN is shown in Figure 2. In the figure, consider Alice who intends to make a payment to Hector, with whom she does not share a payment channel. The naive way to process this transaction would for Alice to open a payment channel with Hector, which involves Alice making an expensive blockchain write for the channel opening. Each payment channel opening costs 2.4 USD for blockchain writes [8, 9, 10]. If Alice intends to send an amount of 1 coin to Hector, it may not be economical for her to open a direct payment channel. Alice can make use of the PCN and make a payment to Hector by forwarding the payment along the path Alice $\rightarrow$ Celia $\rightarrow$ Michael $\rightarrow$ Rajiv $\rightarrow$ Charlie $\rightarrow$ Garcia $\rightarrow$ Hector. This process of using intermediate nodes in a PCN to forward to the payment to the intended destination is called as routing in payment channel networks.
@@
Motivation and timeliness of PCN research: A significant advantage of PCNs is their capability to facilitate micro-payment transactions, with minimum amounts as low as $10^{-7}$ BTC [11]. In contrast, the average transaction cost for a single on-chain transaction on the BTC blockchain is approximately $4.612$ USD as of June 2024 [12]  [^1] Opening a payment channel requires two inputs (for funds from both the parties) and one output (the funds are locked in a single 2-2 multi signature address). Where as making a payment by posting it to the blockchain requires one input (from the party that intends to make the payment) and two outputs (one to the receiver and one to the validator/miner for his fees/reward). This difference makes the transaction cost associated with opening a payment channel cheaper than making a payment using a blockchain write., regardless of the transaction amount. This cost can be avoided by using the off-chain PCNs, which incurs no additional fees. Additionally, transactions on the BTC blockchain take around 2 hours to be confirmed as of June 13, 2024 [5], whereas PCNs can process transactions instantly. An example of a real-world PCN is the Lightning Network ($\mathsf{LN}$) on the Bitcoin blockchain [13], which has a 24-hour trading volume of \$63,200 and a market capitalization of \$7 million [14] as of June 2024. These figures reflect $\mathsf{LN}$’s size and growth.
@@
1) We qualitatively compare the recent work in several aspects of PCNs, viz. pathfinding and routing, virtual channels, state channels, payment channel hubs and rebalancing using several relevant properties (metrics) along with providing a reasoning why these metrics have been chosen for comparison.
@@
2) In this paper, we point out the open problems in all the areas that we survey and we also discuss why solving those problems is a hard research challenge.
@@
Outline: In Section II, we start with describing the concept of pathfinding and routing in PCNs and qualitatively compare work published in that area. In Section III, we describe virtual channels which have been proposed to address issues with multi-hop routing in PCNs, and compare work in this area. In Section III, we describe and compare state channel protocols, which are a more generalized adaptation of virtual channels and can facilitate the execution of arbitrary applications between nodes in the PCN (not just payments). In Section V, we cover payment channel hubs, which are similar to virtual channels and state channels, but facilitate a different use case for payments in PCNs. All of the aforementioned PCN mechanisms consume one common resource: the local balance of a node in a PCN. In Section IV, we discuss rebalancing, which addresses the important function of replenishing channel funds in the PCN. In Section VII, we present the reader with the current research gaps in all of the aforementioned areas and in Section IX, we conclude the paper.
@@
Prior work: Prior works by Khojasteh et al. [15] and Erdin et al. [16] survey the work done only in the area of pathfinding and routing protocols and their privacy aspects in PCNs. Whereas, in this paper we cover the entire spectrum of PCN research: rebalancing, virtual channels and state channels, pathfinding and routing. Apart from this [15, 16] do not provide any information about the open problems in PCNs, which we do in our work. The SoK by Gudgeon et al. [17], surveys several layer-2 protocols, whereas, we focus exclusively on PCNs. Furthermore [17] was published in 2020 and does not cover most of the recent work published in PCNs.
TABLE I: Comparison of Pathfinding and Routing Protocols in PCNs.
@@
Protocols Concurrency Privacy Topology privacy Avoids source routing Decentralized Atomicity Disjoint graphs Fees
MPCN-RP [18] $\checkmark$ $\checkmark$ (flat)
Eckey et al [19] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Vein [20] $\checkmark$ (dynamic)
Auto tune [21] $\checkmark$ (flat)
Kadry et al. [22]
FSTR [23]
SilentWhispers [24] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Blanc [25] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SpeedyMurmurs [26] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Spider [27]
Flash [28] $\checkmark$
Coinexpress [29] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Webflow [30] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Robustpay [31] $\checkmark$ $\checkmark$ (flat)
Robustpay+ [32] $\checkmark$ $\checkmark$ (flat)
RACED [33] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Cheapay [34] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (flat)
Auroch [35] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (dynamic)
Yang et al. [36] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SPRITE [37] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ II Pathfinding and Routing
Motivation: One of the areas in PCNs which has garnered significant attention from the academic community is pathfinding and routing. It is defined as the process of finding a path comprising of several nodes from a sender to a receiver in a PCN along which the payment intended to be sent by the sender can be forwarded to the receiver and forwarding the payment along that path.
Intuitively, it may seem that well-known pathfinding and routing protocols from the networks domain can be easily applied to PCNs. Unfortunately, there are several problems with this.
Even though PCNs are decentralized and distributed peer-peer networks, pathfinding and routing protocols of traditional networks cannot be applied to solve the pathfinding and routing problem in PCNs, because of fundamental differences between intent and the network features: 1) Transfer of data and transfer of money: The intent behind traditional network’s pathfinding and routing protocols is to transfer data between connected devices. In PCNs, the intent to transfer money between users in a decentralized and distributed manner. 2) Change in bandwidth and change in monetary state: When data is transferred between connected devices using network pathfinding and routing protocols, the bandwidth, which is represented by the link weights between the nodes does not change. However, transfer of money between users in a payment channel permanently alters the state of the sender and receiver involved by subtracting and adding money from and to their link weights respectively. 3) Latency and routing fees: The cost of sending data in network pathfinding and routing protocols is measured in terms of latency, which is defined as the time delay experienced in transferring data from one device to another. The delay in data transfer is experienced by the receiver and the cost of the delay is usually paid by the sender. The cost of sending payments in a PCN is measured in terms of money, called as routing fees. This encourages greedy behavior and makes the PCN prone to several attack vectors. Due to these reasons, pathfinding and routing in PCNs is a fundamentally different and a hard research challenge.
@@
The properties in Table I represent the fundamental principles of fiat currency transactions and on-chain payment mechanisms, which we want reflected in off-chain payments. These properties are generally agreed upon in the literature by several works such as [33, 35, 37, 25, 26, 24] as common evaluation metrics for pathfinding and routing protocols in PCNs. Fulfilling these properties while providing efficient pathfinding and routing is a non-trivial challenge, and necessitates the design of new pathfinding and routing protocols. Several elegant pathfinding and routing protocols have been proposed in the literature. In Table I, we present a qualitative comparison of these routing protocols with respect to the properties they achieve.
Ideal properties:
1) Concurrency: Concurrency is achieved when a pathfinding and a routing protocol enables the nodes to forward more than one payment simultaneously.
2) Privacy: Privacy is maintained when a node’s identity is known only to its immediate neighbors and not to the entire network.
3) Topology Privacy: Topology privacy is preserved when no single node has knowledge of the entire network topology.
4) Avoids Source Routing: Source routing is avoided when the sender does not determine the path to the receiver.
5) Decentralization: Decentralization is achieved when there are no centralized or trusted entities responsible for constructing paths for senders.
6) Atomicity: Atomicity is ensured when the payment is routed all the way from the sender to the receiver or the payment is not routed at all.
7) Disjoint Graph Applicability: A pathfinding and routing protocol is considered applicable to disjoint graphs if it functions even when the network graph consists of islands.
8) Fees: Routing fees is the amount a node charges for forwarding the payment to the next node along a path from the sender to receiver. This fees can be charged in two ways. Flat/fixed fees means that the fees charged for routing payments remains the same irrespective of the transaction amount being routed. If the fees charged by a node varies according to the transaction amount, it is referred to as dynamic fees, typically a percentage of the amount being transacted.

### v3

@@ Abstract
Payment Channel Networks (PCNs) have been proposed as an alternative solution to the scalability, throughput, and cost overhead problems associated with blockchain transactions. By facilitating offchain execution of transactions, PCNs significantly reduce the burden on the blockchain, leading to faster transaction processing, reduced transaction fees, and enhanced privacy. Despite these advantages, the current state-of-the-art in PCNs presents a variety of challenges that require further exploration. In this paper, we survey several fundamental aspects of PCNs, such as pathfinding and routing, virtual channels, state channels, payment channel hubs, and rebalancing protocols. We aim to provide the reader with a detailed understanding of the various aspects of PCN research, highlighting important advancements. Additionally, we highlight the various unresolved challenges in this area. Specifically, this paper seeks to answer the following crucial question: What are the various interesting and non-trivial challenges in fundamental infrastructure design leading to efficient transaction processing in PCN research that require immediate attention from the academic and research community? By addressing this question, we aim to identify the most pressing problems and future research directions, and we hope to inspire researchers and practitioners to tackle these challenges to make PCNs more secure and versatile.
@@ I Introduction
Cryptocurrencies and cryptocurrency based transactions have become increasing popular. Currently, the total market value of all cryptocurrencies in use has surpassed 2.5 Trillion USD. The cyrptocurrency market is increasing at a rate of $\approx$ 8.00% every year [1]. This rise in popularity can be attributed to the following reasons: 1) cryptocurrency transactions can be carried out without the presence of a trusted entity. Fiat currency transactions on the other hand, require the presence of a trusted financial organization such as a bank. 2) cryptocurrency transactions do not subject the user to any limits on the number and type of transactions. Fiat currency transactions are limited in their amount and number, and depend on several factors such as the currency, geographical location, etc. [2].
Each transaction posted to the Bitcoin blockchain takes around 7 seconds to be validated [3, 4]. The procedure of validation involves verifying that the transaction posted to the blockchain contains all the required fields and if the signature of the user creating the transaction tuple is valid. Once the validation procedure is successfully completed, the transaction is included in a block that would be mined on the blockchain at some time in the future. The process of mining the block successfully takes $\approx$ 2 hours [5] (as of June $13^{th}$ 2024). This delay in the transactions and blocks getting confirmed is termed as the blockchain scalability problem [6]. In contrast, Visa, a company which globally processes transactions using fiat currencies, processes around 24,000 transactions per second [7]. Due to the delay in transaction processing caused by the blockchain scalability problem, blockchain-based transactions cannot process payments instantaneously.
@@
The idea of a payment channel that exists between two parties can be extended to a number of nodes, creating a network of payment channels, called a payment channel network or PCN. PCNs enable users that are not connected by a direct payment channel to make payments between each other in an off-chain manner. An example PCN is shown in Figure 2. In the figure, consider Alice who intends to make a payment to Hector, with whom she does not share a payment channel. The naïve way to process this transaction would be for Alice to open a payment channel with Hector, which involves Alice making an expensive blockchain write for the channel opening. Each payment channel opening costs 2.4 USD for blockchain writes [8, 9, 10]. If Alice intends to send an amount of 1 coin to Hector, it may not be economical for her to open a direct payment channel. Alice can make use of the PCN and make a payment to Hector by forwarding the payment along the path Alice $\rightarrow$ Celia $\rightarrow$ Michael $\rightarrow$ Rajiv $\rightarrow$ Charlie $\rightarrow$ Garcia $\rightarrow$ Hector. This process of using intermediate nodes in a PCN to forward to the payment to the intended destination is called as routing in payment channel networks.
@@
Motivation and timeliness of PCN research: A significant advantage of PCNs is their capability to facilitate micro-payment transactions, with minimum amounts as low as $10^{-7}$ BTC [11]. In contrast, the average transaction cost for a single on-chain transaction on the BTC blockchain is approximately $4.612$ USD as of June 2024 [12]  [^1] Opening a payment channel requires two inputs (for funds from both the parties) and one output (the funds are locked in a single 2-2 multi signature address). Whereas making a payment by posting it to the blockchain requires one input (from the party that intends to make the payment) and two outputs (one to the receiver and one to the validator/miner for his fees/reward). This difference makes the transaction cost associated with opening a payment channel cheaper than making a payment using a blockchain write., regardless of the transaction amount. This cost can be avoided by using off-chain PCNs, which incur no additional fees. Additionally, transactions on the BTC blockchain take around 2 hours to be confirmed as of June 13, 2024 [5], whereas PCNs can process transactions instantly. An example of a real-world PCN is the Lightning Network ($\mathsf{LN}$) on the Bitcoin blockchain [13], which has a 24-hour trading volume of \$63,200 and a market capitalization of \$7 million [14] as of June 2024. These figures reflect $\mathsf{LN}$’s size and growth.
In this SoK, we do not survey the various types of attacks in PCNs [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 21]. The attacks in PCNs usually have overlaps in their strategy and execution, and most of them currently do not have efficient and fully developed mitigation mechanisms proposed. Our conjecture is that, their countermeasures might also have design overlaps as and when they are proposed. Hence, we believe attacks in PCNs and their countermeasures require their own taxonomy. For this SoK, we have covered papers in various areas of PCNs during the time period of 2019-2024 across Tier-1, Tier-2, Tier-3 security conferences in CS, since the norm for security/privacy research and computer science research in general is peer-reviewed conferences.
@@
1) We qualitatively compare the recent work in several aspects of building PCNs, viz. pathfinding and routing, virtual channels, state channels, payment channel hubs, and rebalancing using several relevant properties (metrics) along with providing a reasoning why these metrics have been chosen for comparison.
@@
2) We point out the open problems in all the areas that we survey and we also discuss why solving those problems is a hard research challenge.
@@
Outline: In Section II, we start with describing the concept of pathfinding and routing in PCNs and qualitatively compare work published in that area. In Section V, we describe virtual channels which have been proposed to address issues with multi-hop routing in PCNs, and compare work in this area. In Section IV, we describe and compare state channel protocols, which are a more generalized adaptation of virtual channels and can facilitate the execution of arbitrary applications between nodes in the PCN (not just payments). In Section V, we cover payment channel hubs, which are similar to virtual channels and state channels, but facilitate a different use case for payments in PCNs. All of the aforementioned PCN mechanisms consume one common resource: the local balance of a node in a PCN. In Section VI, we discuss rebalancing, which addresses the important function of replenishing channel funds in the PCN. In Section VII, we present the reader with the current research gaps in all of the aforementioned areas and also describe why bridging those gaps is hard. In Section VIII, we conclude the paper.
@@
Prior work: Prior works by Khojasteh et al. [42] and Erdin et al. [43] survey the work done only in the area of pathfinding and routing protocols and their privacy aspects in PCNs. Whereas, in this paper we cover the entire spectrum of PCN research: rebalancing, virtual channels, state channels, pathfinding and routing, and tumblers. Apart from this [42, 43] do not provide any information about the open problems in PCNs, which we do in our work. The SoK by Gudgeon et al. [44], surveys several layer-2 protocols, whereas, we focus exclusively on PCNs. Furthermore [44] was published in 2020 and does not cover most of the recent work published in PCNs.
TABLE I: Comparison of Pathfinding and Routing Protocols in PCNs
@@
Protocols Concurrency Privacy Topology privacy Avoids source routing Decentralized Atomicity Disjoint graphs Fees Year
SilentWhispers [45] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2017
SpeedyMurmurs [46] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2018
Coinexpress [47] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2018
Blanc [48] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2019
Robustpay [49] $\checkmark$ $\checkmark$ (flat) 2019
Flash [50] $\checkmark$ 2019
Cheapay [51] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (flat) 2019
Eckey et al [52] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2020
FSTR [53] 2020
Spider [54] 2020
Vein [55] $\checkmark$ (dynamic) 2021
Kadry et al. [56] 2021
Webflow [57] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2021
Robustpay+ [58] $\checkmark$ $\checkmark$ (flat) 2021
MPCN-RP [59] $\checkmark$ $\checkmark$ (flat) 2022
Auto tune [60] $\checkmark$ (flat) 2023
Yang et al. [61] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2023
RACED [62] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2024
Auroch [63] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (dynamic) 2024
SPRITE [64] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2024
@@ II Pathfinding and Routing
Motivation: One of the areas in PCNs that has garnered significant attention from the academic community is pathfinding and routing. Pathfinding is defined as the process of finding a path comprising several nodes from a sender to a receiver in a PCN along which a payment can potentially be forwarded, and routing is the process of actually forwarding the payment along the found path. Intuitively, it may seem that well-known pathfinding and routing protocols from the wired and wireless networks domain can be easily applied to PCNs. Unfortunately, there are several problems with this: 1) Traditional networks focus on the transfer of data, PCNs on the other hand, transfer money in a decentralized manner. 2) Data transfer in traditional peer-peer networks does not alter the bandwidth, whereas money transfer in PCNs alters the monetary state of the nodes involved. 3) Cost in traditional networks is measured in terms of latency, whereas in PCNs, it involves routing fees, leading to greedy behavior among users and makes PCNs vulnerable to various attacks [44].
@@
The properties on the columns in Table I represent the fundamental principles of fiat currency transactions and on-chain payment mechanisms, which we want reflected in off-chain payments. These properties are generally agreed upon in the literature by several works such as [62, 63, 64, 48, 46, 45] as common evaluation metrics for pathfinding and routing protocols in PCNs. Fulfilling these properties while providing efficient pathfinding and routing is a non-trivial challenge, and necessitates the design of new pathfinding and routing protocols. Several elegant pathfinding and routing protocols have been proposed in the literature. In Table I, we present a qualitative comparison of these routing protocols with respect to the properties they achieve. In this paper across all sections, if any prior work has identified a property as ideal or has identified a gap in research, we give an appropriate citation(s). If there is no citation provided, it indicates that the corresponding property/research gap has been identified by us.
Ideal properties: 1) Concurrency: Concurrency is achieved when a pathfinding and a routing protocol enables the nodes to forward more than one payment simultaneously [65]. Importance: At a given instant of time, many users will be using the PCN to make offchain payments. Hence it is important for a routing protocol to support concurrency. 2) Privacy: Privacy is maintained when a node’s real identity is known only to its immediate neighbors and not to the entire network. Importance: Information of a node such as its identity, local balance, connections with other nodes in the network and its transaction history are private and should not be known to anyone else. 3) Topology privacy: Topology privacy is preserved when no single node has knowledge of the entire network topology. Importance: If topology privacy is not preserved, it violates the privacy of every node in the network. Making network topology public can potentially lead an adversary to reconstruct transaction paths, which in turn can lead to an adversary selectively targeting a certain set of nodes. 4) Avoids source routing: Source routing is avoided when the sender does not determine the path to the receiver. Importance: If a sender determines the complete path to the receiver, it means that he has access to the entire network topology. PCNs are highly dense and dynamic in nature. It is practically infeasible for a node to maintain an updated network topology all the time. 5) Decentralization: Decentralization is achieved when there are no centralized, trusted entities responsible for constructing paths for senders. Importance: Cryptocurrency payments made using the blockchain are by nature decentralized, hence routing protocols which facilitate offchain cryptocurrency payments should also be decentralized. 6) Atomicity: Atomicity is ensured when the payment is routed all the way from the sender to the receiver, or the payment is not routed at all. Importance: Atomicity is important since it ensures that honest people do not lose their funds because of malicious behavior by other parties in the system. 7) Disjoint graphs: A pathfinding and routing protocol is considered applicable to disjoint graphs if it functions even when the network graph consists of islands. Importance: PCNs often comprise of islands which only have a couple of nodes. These islands are often disconnected from other dense parts of the PCN. A routing protocol should be able to facilitate transactions between any pair of nodes irrespective of their location. 8) Fees: Routing fees is the amount a node charges for forwarding the payment to the next node along a path from the sender to receiver. This fees can be charged in two ways. Flat/fixed fees means that the fees charged for routing payments remains the same irrespective of the transaction amount being routed. If the fees charged by a node varies according to the transaction amount, it is referred to as dynamic fees, typically a percentage of the amount. Importance: Every node along a payment path aids the sender in transaction processing by forwarding the payment to the next node along the path. The nodes need to be paid in the form of routing fees for their service.
As illustrated in Table I, routing protocols for PCNs have evolved significantly over the years. The two most significant advancements are taking routing fees into consideration and providing support for privacy. For instance, $\mathsf{LN}$ provides users with two sets of keys, a long-term keypair and an alias (a temporary identity) helping to conceal their identities and ensure privacy. Despite these developments, two overarching research problems remain that require attention. We discuss them in detail in Section VII.